OpenStudy (anonymous):

show that sin(t)+cos(t)i=cis(pi/2-t). how do i do this? related to complementary angles of cos and sine ?

OpenStudy (anonymous):

\[\sin (\frac{\pi}{2}-t)=\cos(t)\]

OpenStudy (anonymous):

\[\cos (\frac{\pi}{2}-t)=\sin(t)\]

OpenStudy (anonymous):

these are the only substitutions that you will need.